# Enhancing Large Language Model-based Speech Recognition by Contextualization for Rare and Ambiguous Words

### v1

@@ 2.1 Model Design
We adopt an in-house 100B-parameter LLM pre-trained on Japanese and English datasets from scratch, referred to as PLaMo-100B [^1] The details of this model will be published. The blog post briefly explaining this model and datasets is available at https://tech.preferred.jp/ja/blog/plamo-100b in Japanese., as a decoder, and the encoder of pre-trained Whisper large-v3222We used the checkpoint at https://huggingface.co/openai/whisper-large-v3. as an audio encoder, as shown in Figure 1. We also construct an LLM-based ASR system using Swallow-7B (Fujii et al., 2024) as a decoder that performed continual pre-training on LLaMA-2 7B (Touvron et al., 2023) on Japanese and English datasets, where Japanese was dominant.
@@ 5 Conclusion
Kento Nozawa implemented a training and evaluation pipeline, conducted numerical experiments, and drafted the manuscript. Takashi Masuko constructed all datasets and extracted their keywords, implemented keywords error evaluation, and wrote the manuscript. Toru Taniguchi prototyped and supported training and evaluation pipelines, revised the manuscript, and managed this project.
@@ Acknowledgments
We would like to thank (ex-) PFN/PFE members, especially Daiki Higurashi, Linsho Kaku, Toshiki Kataoka, Hiroaki Mikami, Shintarou Okada, Daisuke Okanohara, Shuji Suzuki, Seiya Tokui, Zijian Xu, and Shoichiro Yamaguchi, for their helpful discussions and/or implementation support; Toshihiko Yanase, Yuta Hirokawa, and Toru Komatsu for infrastructure-related support.

### v2

@@ 2.1 Model Design
We adopt an in-house 100B-parameter LLM pre-trained on Japanese and English datasets from scratch, referred to as PLaMo-100B (Preferred Elements et al., 2024), as a decoder, and the encoder of pre-trained Whisper large-v3111We used the checkpoint at https://huggingface.co/openai/whisper-large-v3. as an audio encoder, as shown in Figure 1. We also construct an LLM-based ASR system using Swallow-7B (Fujii et al., 2024) as a decoder that performed continual pre-training on LLaMA-2 7B (Touvron et al., 2023) on Japanese and English datasets, where Japanese was dominant.
@@ 5 Conclusion
Kento Nozawa implemented the training and evaluation pipeline, conducted numerical experiments, and drafted the manuscript. Takashi Masuko constructed all datasets and extracted their keywords, implemented keywords error evaluation, and wrote the manuscript. Toru Taniguchi prototyped and supported the training and evaluation pipeline, revised the manuscript, and managed this project.
@@ Acknowledgments
We would like to thank (ex-) PFN/PFE members, especially Daiki Higurashi, Linsho Kaku, Toshiki Kataoka, Hiroaki Mikami, Shintarou Okada, Daisuke Okanohara, Shuji Suzuki, Daisuke Tanaka, Seiya Tokui, Zijian Xu, and Shoichiro Yamaguchi, for their helpful discussions and/or implementation support; Toshihiko Yanase, Yuta Hirokawa, and Toru Komatsu for infrastructure-related support.